OpenStudy (anonymous):

use the binomial theorem to find the coefficient of x^4 in the expansion of (3x-1)^10

OpenStudy (aum):

\[ \large \text{k-th term of the Binomial Series: } \\ \large (a + b)^{n} = \left(\begin{matrix}n \\ k\end{matrix}\right)a ^{n - k}b ^{k} ~~~~~~~ \text{where }\left(\begin{matrix}n \\ k\end{matrix}\right) = \frac{ n! }{ (n-k)! k! } \\ \large \\ \large \text{Substitute: }~~a = 3x;~~ b = -1; ~~ n = 10; ~~ k = 6 \]